Shift summary Day 10 November 1999.

Nickolas Solomey (solomey@hep.uchicago.edu)
Wed, 10 Nov 1999 15:28:22 -0600

Shift Summary for Day shift, 10/Nov/99

On Shift: Roland and Nick

Hours of beam available at all intensities: 7.45
Hours of beam available at nominal intensity: 7.35
Hours of E799 data (special runs not included): 7.35 hr
Hours of special runs (laser scans, muon runs...): 0 hr

Accounting of lost time: 45 min for down tevatron due to
quench plus odd spills here and there.

========================================================================
Runs ended this shift: None
Run 14938 running whole shift

Tapes brought to FCC this shift: None
(including tapes for online split)

-------------------------------------------------------------------------
Accesses during shift: None

========================================================================
Overview of shift: Mostly data taking except for one short tev
quench when we were off.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14938 76 07:00 Shift swap.
122 08:05 Trigger hung, no verbal announcement,
did kfrend anyway and all was fine.
127 Request to mount split tape UPF032
on ktev4 drive tps3d2.
162 Ped problem, channel 495 and 499,
reset pipeline with dbc and all came
back to normal. (Julie listed as
CsI problem contact was paged but
did not call back so we did pipeline
reset and it fixed the problem.)
168 Seems to be good now.
170 Accelertor missed a spill, maybe I
should call and complain!
177 Trigger hung in error state, did
kfrend but would not start. Cycled
power on downstairs remote unit
and kfrend (Jason L suggestion) then
triggers were back by spill 180.
181 Triggers seem good.
207 09:56 Acc. Abort, half a spill missing.
208 Beam off.
212 10:03 Beam returns, no MCR message.
227 Half intensity spill
231 10:30 Split requested tape UPX010 mounted
on ktev3, still at half intensity.
234 Spill return to nominal 9e12 intensity.
380 13:52 On split requested tap UPT010 mounted
on ktev2.
382 No beam, Tevatron quench!
410 14:30 Some beam returns, 1e12.
413 beam, 8.6e12, good spill structure!
However, Jason claims he has seen
better; as for this shift it is the
best spill structure thus far today.
447 15:20 No beam this spill.
452 Beam gone.
455 15:30 Shift ends, run is ongoing.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey